# SAM Decoding: Speculative Decoding via Suffix Automaton

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing by unifying tasks into text generation, yet their large parameter sizes and autoregressive nature limit inference speed. SAM-Decoding addresses this by introducing a novel retrieval-based speculative decoding method that uses a suffix automaton for efficient and accurate draft generation. Unlike n-gram matching used by the existing method, SAM-Decoding finds the exact longest suffix match in generating text and text corpuss, achieving an average time complexity of $O(1)$ per generation step. Meanwhile, based on the exact match length, it is designed as an approach that can be combined with existing methods, allowing SAM-Decoding to adaptively select a draft generation strategy based on the matching length, thus increasing the inference speed of the LLM. When combined with Token Recycling, evaluations show SAM-Decoding outperforms existing model-free methods, achieving a speedup of 2.27×2.27$\times$2.27 × over autoregressive decoding on Spec-Bench. When combined with EAGLE2, it reaches a speedup of 2.49×2.49$\times$2.49 ×, surpassing all current approaches. Our code is available at ours repository.
@@
Yuxuan Hu1,2, Ke Wang1,2, Xiaokang Zhang1,2, Fanjin Zhang4 Cuiping Li1,3Hong Chen1,3Jing Zhang1,3thanks: Corresponding author. 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3Engineering Research Center of Database and Business Intelligence, Beijing, China 4Knowledge Engineering Group, Tsinghua University, Beijing, China
@@ 1 Introduction
The Transformer-based Large Language Model Brown et al. (2020); Dubey et al. (2024); Yang et al. (2024) has become a foundation model in the field of natural language processing. Through the large language model (LLM), all natural language processing tasks are unified as text generation tasks, which greatly improves the convenience of solving natural language processing tasks. While LLM has shown good performance in various tasks and is easy to use, the large-scale parameters and text generation-based paradigm contained in LLM have brought challenges to the practical deployment, especially in terms of the inference speed of the LLM. LLM generates tokens step by step in an autoregressive manner. At each round of generation, LLM generates one new token based on the tokens that has been generated. In addition, for the GPUs used for deploy LLM, this process requires reading the huge number of parameters contained in the LLM from the High-Bandwise Memory (HBM) to the on-chip memory, but these parameters only operate on a small number of parameters corresponding to the new token, resulting in the inability to fully utilize the parallel capabilities of the modern GPU, thereby limiting the inference speed of the LLM.
To address the above problem, one effective strategy is to enhance the arithmetic intensity of the large language model (LLM) generation process, which involves reducing the number of generation rounds while increasing the number of tokens generated in each round. Building upon this concept, the method of speculative decoding has been introduced and has spurred a series of follow-up studies Leviathan et al. (2023); Miao et al. (2024); Cai et al. (2024); Fu et al. (2024). Rather than generating a single token per step, these approaches employ a draft model to predict token sequence as draft. The predicted sequence are then verified against the LLM, and only the accurate tokens are adopted.
Most speculative decoding techniques employ various neural networks, such as smaller language models (LLMs) or embedding layers, as draft models, which we categorize as model-based methods. Although model-based speculative methods perform well on all kinds of tasks, when we consider some special tasks such as summarization, rewriting, there is a simpler type of speculative decoding methods that exhibit superior performance, which we call retrieval-based speculative decoding methods. These methods involve retrieving potential subsequent tokens from previously generated content or a text database during each generation round. Unlike model-based approaches, retrieval-based methods do not necessitate the training of draft models or the use of GPUs for inference.
Figure 1: Overview of SAM-Decoding’s workflow. SAM-Decoding includes two automaton: static SAM and dynamic SAM. Among them, the static SAM is built based on a given text corpus, while the dynamic SAM is built based on generating text and gradually expanded during the generation process. When drafting, the suffix automaton matches the suffixes of the generating text and retrieves the draft from the text corpus and the generating text respectively according to the matching position. Meanwhile, we also combined an auxiliary speculative decoding algorithm (Aux Method) to deal with scenarios where retrieval is not applicable. we select the draft to be used from the three candidate drafts based on the match length. After that, the draft will be verified by LLM to obtain accepted tokens. Using these accepted tokens, we finally extend the dynamic SAM and generating text for the next round of generation.
After examining the current retrieval-based methods, we identify several limitations. Firstly, these approaches typically rely on a single retrieval source. For instance, PLD Saxena (2023) focuses solely on prompts, whereas REST He et al. (2024) is limited to text corpus. Secondly, there are limitations in the retrieval techniques they adopt; PLD relies on n-gram matching, which is fast for context matching but cannot be extended to a wider range of textual repositories due to its poor theoretical complexity, while REST is based on suffix arrays, which has better complexity but cannot be applied to continually generating text. Lastly, not every step in the generation is suitable for the retrieval method; in a few generation steps, a large number of tokens can be accepted based on PLD and REST, but in the rest of the generation steps there is no acceleration.
To address the limitations of the retrieval-based approach mentioned earlier, this paper introduces SAM-Decoding, an innovative retrieval-based speculative decoding technique. Unlike existing approaches that rely on n-gram matching, SAM-Decoding employs a suffix automaton (SAM) to solve the exact longest suffix match problem for the generating text within both the historical generated text and the text corpus. This process yields more accurate match positions and exact match lengths compared to n-gram matching. Moreover, during the text generation phase, the suffix automaton effectively captures the relationships between adjacent suffixes, ensuring that the average time complexity for identifying the longest suffix match at each step remains at $O(1)$. This represents a significant improvement over the theoretical computational efficiency of n-gram matching. Meanwhile, the length of the longest suffix match serves as an indicator of the alignment between the generating text and retrieval sources. By exploiting this metric, our approach can integrate existing speculative decoding methods as an auxiliary and dynamically decide whether to generate a draft based on the matched content, or utilize an auxiliary decoding method. This adaptability leads to a notable increase in the overall speed of the text generation process.
@@
Specifically, SAM-Decoding first constructs a static suffix automaton for the existing text corpus. For each request, SAM-Decoding then builds a dynamic suffix automaton based on the input prompt. The suffix automaton contains a series of nodes, each of which corresponds to several substrings in the context (text corpus, prompt and generating text). We record the earliest corresponding position of each node in the text. This information facilitates the generation of drafts by directly mapping nodes in the automaton to indices in the context. We also maintain the corresponding nodes of the generating text in both the static and dynamic automaton. Therefore, in each round of generation, we can directly retrieve the draft from the context according to the position recorded in the node, and filter the draft according to the matching length. After generation, as new tokens are generated, for the static automaton, we transition nodes of SAM based on new tokens, and for the dynamic automaton, we first expand the structure to accommodate the new tokens before transitioning nodes of SAM accordingly.
Extensive evaluations across a range of tasks indicate that the SAM-Decoding can achieve results that are competitive with existing state-of-the-art (SOTA) approaches in all tasks. Particularly in tasks where retrieval-based methods are applicable, SAM-Decoding outperforms all existing methods. On Spec-Bench, the latest benchmark for speculative decoding, SAM-Decoding combined with Token Recycling Luo et al. (2024) achieves a speedup of 2.26 times compared to autoregressive decoding, outperforming all model-free baselines. Additionally, when SAM-Decoding is combined with EAGLE2 Li et al. (2024a), the SOTA model-based speculative decoding method, it achieves a speedup of 2.49 times compared to autoregressive decoding, outperforming all existing model-based and model-free methods.
@@ 2.1 Suffix Automaton
A suffix automaton contains a series of nodes and two types of state transfer edges, extension edges and suffix link edges. A node in the automaton represents a state and corresponds to all substrings that have the same ending position in the string. Meanwhile, extension edges are standard edges that represent a possible extension of the current substring by appending a new character, while suffix link edges create a path that allows the automaton to quickly jump to states representing shorter suffixes of the current substring.
@@ 2.2 Speculative Decoding
Given the model input $x=(x_{1},x_{2},\ldots,x_{t})$, an LLM generates a new token $x_{t+1}$ at each generation step autoregressively. The key idea of speculative decoding is to utilize a lightweight draft model to generate multiple candidate tokens quickly, i.e., $x_{\text{draft}}=(x_{t+1},x_{t+2},\ldots,x_{t+n})$, and then the target LLM simultaneously evaluates these candidates and accept those aligned with the output distribution of the LLM, i.e., $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$.
In the above, we assume that the draft is a sequence of tokens. Recent works proposed to generate a candidate token tree via a tree mask in the attention module to make the target LLM simultaneously evaluate multiple branches of this token tree, thereby increasing the acceptance length of the draft model.
@@ 3 SAM-Decoding
To tackle the first two limitations, SAM-Decoding leverages suffix automaton, which significantly enhances the coverage of retrieved corpus and the efficiency of the retrieval process while allowing for more flexible matching lengths. In what follows, we detail how SAM-Decoding can be integrated with both model-free and model-based methods. By utilizing the precise matching information provided by the suffix automaton, our method not only overcomes the third limitation but also ensures consistent performance improvements across a wide range of tasks.
@@ 3.1 Suffix Automaton Construction
SAM-Decoding facilitates rapid draft generation by leveraging suffix automaton. Unlike traditional methods that rely on a single retrieval source, our approach utilizes both a text corpus and the generating text itself as retrieval sources. To accommodate these dual sources, we construct two distinct types of suffix automaton: a static suffix automaton and a dynamic suffix automaton. For the text corpus, we pre-build a static suffix automaton offline. During the inference phase, state matching is performed exclusively on this static automaton. Conversely, for the generating text, we create a dynamic suffix automaton that incrementally expands as the text generation progresses. State matching is also carry out concurrently with this expansion.
A suffix automaton can be constructed in linear time using Blumer’s algorithm Blumer et al. (1984). Since the suffix automaton is designed for a single reference string, when dealing with multiple strings in the text corpus, we concatenate them into a single long string using special symbols, such as an End-of-Sentence (EOS) token. We then construct a static suffix automaton for this concatenated string.
At the same time, we also made several modifications to the suffix automaton to better suit draft generation. Specifically, at each generation step, the generating text corresponds to a node in the automaton, representing the longest suffix match. We need to identify the matching position in either the text corpus or the generating text. Based on this position, we can use the subsequent tokens as potential drafts. To achieve this, at every node of the suffix automaton, we record the earliest position of all substrings corresponding to that node in the reference string. We define this position as min_endpos. By using the min_endpos, we can directly map node to the position in the reference string. The construction process of the suffix automaton is detailed in Appendix A.1.
@@ 3.2 Drafting with Suffix Automaton
After constructing the suffix automaton, it can be utilized as a draft model to generate draft at each step of the generation process. Specifically, let the current input to the LLM be denoted as $x=(x_{1},x_{2},\ldots,x_{t})$, with the predicted next token being $x_{t+1}$. We define the suffix automaton as $S$ and its associated reference text as $T$. The state within the suffix automaton corresponding to the sequence $x$ is denoted as $s_{t}$. The transition to the next state is performed based on the newly generated token $x_{t+1}$ and the current state $s_{t}$:
@@
Subsequently, we extract $n$ consecutive tokens from the reference text $T$ to form a draft, using the min_endpos value stored in node corresponding to state $s_{t+1}$. If the min_endpos value in $s_{t+1}$ is $p_{t+1}$, then the draft $d$ is defined as:
$\text{\rm d}=T[p_{t+1}:p_{t+1}+n],$
where $d$ represents the generated draft and $n$ denotes the length of the draft.
In practical use, we utilize two types of automaton: a static automaton for the text corpus and a dynamic automaton for generating text. To distinguish between the two, we explicitly record the attribute $l_{t+1}$ for the state $s_{t+1}$. This attribute represents the longest suffix length matched by the generating text in the automaton. This helps us determine whether the draft should be generated using the static suffix automaton or the dynamic suffix automaton. Our experimental findings indicate that drafts generated from the dynamic automaton often outperform those from the static text corpus. Consequently, we prioritize drafts from the dynamic automaton. Specifically, let $l_{1}$ and $l_{2}$ represent the matching lengths of the static automaton and the dynamic automaton, respectively. We use the draft provided by the static automaton only if $l_{1}>l_{2}+l_{\text{bias}}$; otherwise, we use the draft provided by the dynamic automaton. Here, $l_{\text{bias}}$ is a predefined constant greater than 0.
We illustrate the state transfer process of the suffix automaton in Algorithm 1. Using amortized analysis, we can prove that the average complexity of state transfer is $O(1)$, while the worst-case time complexity is $O(L)$, where $L$ represents the length of the generating text. The detailed proof is provided in Appendix A.2. Reviewing existing methods, PLD uses $n$-grams and performs a brute-force search for matches in the prompt, resulting in a time complexity of $O(n^{2}L)$. REST also employs $n$-grams but searches for matches in the text using suffix arrays, leading to a time complexity of $O(n^{2}\log L)$. Here, $n$ represents the predefined maximum matching length, and $L$ represents the length of the prompt and the total text in the corpus, respectively. In comparison, our proposed model (SAM-Decoding) has a lower time complexity. Additionally, SAM-Decoding can find the exact longest suffix match without any limit on the matching length. This means that SAM-Decoding can generate drafts both faster and more accurately.
@@
        $s=s.\text{link}$
        $l=s.\text{length}$
@@
     return $s$, $l$

### v3

@@ Abstract
Speculative decoding (SD) has been demonstrated as an effective technique for lossless LLM inference acceleration. Retrieval-based SD methods, one kind of model-free method, have yielded promising speedup, but they often rely on incomplete retrieval resources, inefficient retrieval methods, and are constrained to certain domains. This paper presents a novel retrieval-based speculative decoding method that adapts suffix automaton (SAM) for efficient and accurate draft generation by utilizing common text corpus and dynamic text sequence. Unlike existing $n$-gram matching methods, SAM-Decoding finds the exact longest suffix match, achieving an average time complexity of O(1) per generation step of SAM update and suffix retrieval. It can also integrate with existing methods, adaptively selecting a draft generation strategy based on match length to generalize to broader domains. Extensive experiments on Spec-Bench show that our method is $18\%+$ faster than other retrieval-based SD methods. Additionally, when combined with advanced EAGLE-2, it provides an additional speedup of $3.28\%$$11.13\%$ across various-sized LLM backbones. Our code is available at our repository.
@@
Yuxuan Hu1,2, Ke Wang1,2, Xiaokang Zhang1,2, Fanjin Zhang4 Cuiping Li1,3Hong Chen1,3Jing Zhang1,3thanks: Corresponding author. zhang-jing@ruc.edu.cn 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3Engineering Research Center of Database and Business Intelligence, Beijing, China 4Knowledge Engineering Group, Tsinghua University, Beijing, China
@@ 1 Introduction
The Transformer-based Large Language Models (LLMs) Brown et al. (2020); Dubey et al. (2024); Yang et al. (2024) have demonstrated remarkable abilities and are extensively adopted in numerous domains. The scaling law drives LLMs to become deeper, reaching hundreds of billions of parameters, which makes them inefficient for generating text in a token-by-token autoregressive manner. Speculative decoding methods Leviathan et al. (2023); Cai et al. (2024) seek to tackle this problem by quickly generating multiple draft tokens and subsequently concurrently verifying them with LLMs. These methods can decrease inference latency substantially while maintaining decoding accuracy.
Speculative methods can be categorized into model-based and model-free methods. Model-based methods need to carefully choose and train one or more small-sized draft models. For example, Medusa Cai et al. (2024) utilizes multiple decoding heads to generate multiple future tokens while EAGLE-2 Li et al. (2024a) leverages shallow Transformer layers to predict the next last hidden states and corresponding decoding tokens. Although these methods achieve impressive speedup, they often fail to generate long draft tokens due to drafting overhead or decaying prediction accuracy. Retrieval-based speculative decoding methods, a major type of model-free methods, aim to remedy this issue by generating draft tokens from text corpus or current text sequence.
Figure 1: Throughput of Vicuna-7B, Vicuna-13B, Vicuna-33B on MT-Bench with A6000 GPU using PLD, Token  Recycling Luo et al. (2024), EAGLE-2, and SAM-Decoding, where PLD is the SOTA retrieval-based SD baseline.
However, current retrieval-based methods have notable limitations. Firstly, diverse retrieval sources contribute to the efficiency of retrieval-based SD methods, but existing methods typically rely on a single retrieval source: PLD Saxena (2023) focuses on current text while REST He et al. (2024) uses a text corpus. Secondly, the retrieval techniques they use have efficiency limitations. PLD finds $n$-gram matching from current text sequence, but it has poor theoretical computational complexity and limited applicability to larger text corpus. REST uses suffixed arrays, which provides better complexity than PLD, but still not optimal complexity. Thirdly, retrieval-based methods are suitable for specialized domains (e.g., summarization and RAG), which are unable to bring a noticeable acceleration in other domains.
To address limitations in previous retrieval-based methods, this paper introduces SAM-Decoding, an innovative speculative decoding technique based on suffix automaton. (1) To enhance the coverage of the retrieved corpus, we utilize the common text corpus and the current text sequence as retrieved sources. (2) To improve the retrieval efficiency and accuracy, we adapt a suffix automaton (SAM) to solve the longest suffix match problem, which yields more accurate match positions and exact match length compared to $n$-gram matching. As for retrieval efficiency, the average time complexity of SAM update and suffix retrieval is $O(1)$ by capturing relationships between adjacent suffixes. (3) To generalize our method, assuming the matching length of the longest suffix implying the quality of retrieval draft tokens, our method can be integrated with other types of speculative decoding methods, enabling more efficient text generation by deciding whether to adopt auxiliary decoding techniques.
@@
Specifically, SAM-Decoding creates both a static suffix automaton for the text corpus and a dynamic suffix automaton for the current text sequence. The nodes of suffix automaton represent substrings in the text corpus or current sequence. The earliest position of each substring is recorded in each node. During generation, we can directly retrieve and filter drafts from the context using the matching positions and longest suffixes’ matching length. After each generation step, the automaton is updated: static automaton nodes transition based on new tokens, while the dynamic automaton first expands its structure before node transitions.
Extensive evaluations demonstrate the competitive performance of our method across tasks. On Spec-Bench, SAM-Decoding achieves $18\%+$ faster than previous retrieval-based speculative decoding methods (e.g., PLD, REST, etc.). SAM-Decoding further achieves speedups of up to 1.3×1.3$\times$1.3 × over alternative baselines on the code-generation benchmark like HumanEval. When combined with EAGLE-2 Li et al. (2024a), as shown in Figure 1, our method outperforms the state-of-the-art, delivering an additional $3.28\%$$11.13\%$ speedup on MT-Bench w.r.t. various LLM backbones.
@@ 2 Background
Figure 2: The suffix automaton corresponding to the string “ABCBC”.
@@ 2.1 Suffix Automaton
As shown in Figure 2, a suffix automaton contains a series of nodes and two types of state transfer edges, extension edges (next) and suffix link edges (link). A node in the automaton represents a state and corresponds to all substrings that have the same ending position in the string. Meanwhile, extension edges are standard edges that represent a possible extension of the current substring by appending a new character, while suffix link edges create a path that allows the automaton to quickly jump to states representing shorter suffixes of the current substring.
@@
Figure 3: Overview of SAM-Decoding’s workflow. In each round of generation, the suffix automaton matches the suffixes of the generating text and retrieves the draft from the text corpus and the generated text respectively according to the matching position. Our method can be combined with an auxiliary SD algorithm (Auxiliary) to deal with the scenarios where the retrieval is not applicable. We select the best draft from the three candidate drafts based on the match length, and then the drafts are verified by the LLM for accepted tokens. Using these accepted tokens, we finally extend the dynamic SAM and generate text for the next round of generation.
@@ 2.2 Speculative Decoding
Given the model input $x=(x_{1},x_{2},\ldots,x_{t})$, an LLM generates a new token $x_{t+1}$ at each generation step autoregressively. The key idea of speculative decoding is to utilize a lightweight draft model to generate multiple candidate tokens quickly, i.e., $x_{\text{draft}}=(x_{t+1},x_{t+2},\ldots,x_{t+n})$, and then the target LLM simultaneously evaluates these candidates and accept those aligned with the output distribution of the LLM, i.e., $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$, where $n$ and $m$ denote the size of the draft and the number of accepted tokens.
In the above, we assume that the draft is a sequence of tokens. Recent works proposed to verify a candidate token tree via a tree mask in the attention module to make the target LLM simultaneously evaluate multiple branches of this token tree, thereby increasing the acceptance length of the draft model.
@@ 3 SAM-Decoding
To tackle the first two limitations, SAM-Decoding leverages suffix automaton on diverse text sources, which significantly enhances the coverage of retrieved corpus and the efficiency of the retrieval process while allowing for flexible matching lengths. In what follows, we detail how SAM-Decoding can be integrated with both model-free and model-based methods. By utilizing the precise matching information provided by the suffix automaton, our method not only overcomes the third limitation but also ensures consistent performance improvements across a wide range of tasks. The workflow of SAM-Decoding is shown in Figure 3.
@@ 3.1 Suffix Automaton Construction
To cover comprehensive retrieval sources, SAM-Decoding builds suffix automaton (SAM) by utilizing common text corpus and the current text sequence (including user prompts and already generated tokens). Thus, we construct two types of suffix automaton: a static suffix automaton and a dynamic suffix automaton. For the text corpus, we pre-build a static suffix automaton offline, which is used for state matching during inference. For the current text sequence, we create and expand a dynamic suffix automaton incrementally as generation progresses and performing state matching concurrently.
A suffix automaton can be constructed in linear time using Blumer’s algorithm Blumer et al. (1984). Since the suffix automaton is designed for a single reference string, static suffix automation can not be directly built using a text corpus. To this end, we concatenate multiple strings in the corpus by using special symbols like an End-of-Sentence (EOS) token. We then construct a static suffix automaton for this concatenated string.
We have modified the suffix automaton for better draft generation. During each generation step, the current generated text sequence corresponds to a node in the automaton, representing the longest suffix match. At each node of the suffix automaton, we record the earliest position of all substrings corresponding to that node in the reference string, termed as min_endpos, which allows us to efficiently locate the previous ending position of the matched longest suffix. Hereafter, the subsequent tokens after the matched suffix can be regarded as potential drafts. The construction process of the suffix automaton is detailed in Appendix A.1.
For the static suffix automaton, based on the frequency of occurrence of different substrings, we additionally compute the top-k successor states (topk_succ) of each state, and subsequently use them to construct more complex tree drafts. Although computing the successor states requires significant computation, this can be done offline, eliminating the need to account for this time overhead in real-time processing.
@@ 3.2 Drafting with Suffix Automaton
We illustrate how to generate draft tokens efficiently based on the built suffix automaton. Let $S$ denote the suffix automaton, $T$ denote its associated reference text, and $x=(x_{1},x_{2},\ldots,x_{t})$ denote the current text sequence. The state within the suffix automaton corresponding to the sequence $x$ is denoted as $s_{t}$. In each round of generation, the transition to the next state is performed based on the newly generated token $x_{t+1}$ and the current state $s_{t}$:
@@
For dynamic suffix automaton, we extract $n$ consecutive tokens from the reference text $T$ to form a draft, using the min_endpos value stored in the node corresponding to state $s_{t+1}$, termed as $p_{t+1}$. Then the draft $d_{t+1}$ is defined as:
$d_{t+1}=T[p_{t+1}+1:p_{t+1}+n],$
where $d_{t+1}$ represents the generated draft and $n$ denotes the length of the draft.
For static suffix automaton, we construct a tree-structured draft by Prim’s algorithm based on top-k successors, as detailed in Appendix A.2,
$d_{t+1}=\text{Prim}(S,s_{t+1},x_{t}).$
In practical use, we track the longest-matched suffix length (denoted as $l$) to determine whether to use the static suffix automaton or the dynamic suffix automaton. Specifically, let $l_{1}$ and $l_{2}$ be the matching lengths of the static and dynamic automata, respectively. Our experimental findings indicate that drafts generated from the dynamic automaton often outperform those from the static text corpus. Consequently, we prioritize drafts from the dynamic automaton. We use the draft from the static automaton only if $l_{1}>l_{2}+l_{\text{bias}}$, where $l_{\text{bias}}$ is a predefined constant.
The complete state transfer process of the suffix automaton is shown in Algorithm 1. Using amortized analysis, we can prove that the average complexity of state transfer is $O(1)$, with a worst-case time complexity of $O(L)$, where $L$ is the length of the current generated text (C.f. proof in Appendix A.3). Existing methods like PLD uses a brute-force search for $n$-gram matches, resulting in a time complexity of $O(n^{2}L)$. REST also employs n-grams but searches using suffix arrays, leading to a time complexity of $O(n^{2}\log L)$. Here, $n$ is the predefined maximum matching length, and $L$ is the length of the current text or the concatenated texts in the corpus. In contrast, our proposed SAM-Decoding model has a lower time complexity and can find the exact longest suffix match without any limit on matching length, making it faster and more accurate for draft generation.
@@
        $s,\,l=s.\text{link},\,s.\text{link}.\text{length}$
@@
     Output: next state $s$, next matching length $l$